OpenStudy (anonymous):

A math club had 7 members in its first year. It had 22 members in its fourth year and 32 members in its sixth year. Assuming a constant rate of increase, how many members will the math club have in its thirteenth year?

OpenStudy (anonymous):

in 3 years 15 members came into the club in 2 years 10 members came into the club

OpenStudy (anonymous):

67

OpenStudy (anonymous):

(22-7)/(4-1) = 5 (so every year the number increases by 5) 13th year = (5X12)+7=67

OpenStudy (anonymous):

there are 67 members

OpenStudy (anonymous):

thanks guys!